Suppose that two firms emit the same chemical pollutant when producing good X. The marginal cost...

  1. Suppose that two firms emit the same chemical pollutant when producing good X. The marginal cost of reducing pollution for each firm is as follows: MAC1 = 30 + 5Z1 and

MAC2 = 15 + 20Z2, where Z is the amount of pollution reduction. The marginal social benefit (MSB) of total pollution reduction is MSB = 500 – 5ZT, where ZT = Z1 + Z2.

  1. How much pollution reduction does each firm should contribute in order to achieve the allocative efficient level of abatement? Show all your work.

  1. What is the total allocative efficient level of pollution reduction? Show all your work.
Answer #1

MAC 260 / u0 30+52; 19AC, ATAC,+MAC, MSB -93725 120 12 = 41146 5 10 15 20 25 30 35 40 w 50 55 60st oooo MACAMACA : MSB. -,- 232.95 V 129 5 10 15 20 25 30 35 nous 505560L e condition for allocative efficiency is MAC, = 30+521 MAC2 = 15+2022 MAC, MACZ = MSB. first equate MAC, & MAC2 30+52,315+2ZT. = 2, + Z. - 41.4444+1). U -52-55 MSB: 500 -5(27722). 500-5(52-553 - 500 - 262.75 2 237.25 )